# Screened Thin-Target Bremsstrahlung with Partially-Ionized High-Z Species

### v1

@@ III.2 Screened Cross Section
$\frac{d^{3}\sigma_{Z_{s^{\prime},s}}^{\text{MY}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}=\alpha\left(\frac{r_{e}}{2\pi}\right)^{2}Z_{s}^{2}\times$ (35)
$\left(1-\left(1-q_{Z_{s^{\prime},s}}\right)\sum_{i}\frac{\overline{A}_{s^{\prime},s,i}}{1+\left(\bar{q}\bar{a}_{Z_{s^{\prime},s,i}}/2\right)^{2}}\right)^{2}\frac{\bar{p}}{\bar{k}\bar{p}_{0}\bar{q}^{4}}A^{\text{BH}},$ (36)
where the term $q_{Z_{s^{\prime},s}}:=Z_{s^{\prime},s}/Z_{s}$, ranging between 0 and $1$, characterizes the level of ionization, 0 being the neutral atom. Equation (35) may then be rewritten in the form
$\frac{d^{3}\sigma_{Z_{s^{\prime},s}}^{\text{MY}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}$ $=\left(\sum_{i}\overline{A}_{s^{\prime},s,i}\frac{\bar{q}^{2}+q_{Z_{s^{\prime},s}}\left(2/\bar{a}_{Z_{s^{\prime},s,i}}\right)^{2}}{\left(2/\bar{a}_{Z_{s^{\prime},s,i}}\right)^{2}+\bar{q}^{2}}\right)^{2}$ (37)
$\times\frac{d^{3}\sigma_{Z_{s}}^{\text{BH}}}{d\bar{k}d\Omega_{k}d\Omega_{p}},$ (38)
and defining $\overline{b}_{s^{\prime},s,i}:=2/\bar{a}_{Z_{s^{\prime},s,i}}$, a simple and compact formulation may be obtained
$\frac{d^{3}\sigma_{Z_{s^{\prime},s}}^{\text{MY}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}=\left(\sum_{i}\overline{A}_{s^{\prime},s,i}\frac{\bar{q}^{2}+q_{Z_{s^{\prime},s}}\overline{b}_{s^{\prime},s,i}^{2}}{\bar{q}^{2}+\overline{b}_{s^{\prime},s,i}^{2}}\right)^{2}\frac{d^{3}\sigma_{Z_{s}}^{\text{BH}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}.$ (39)
From now on, we omit the BH superscript on the Bethe–Heitler TDCS for the sake of readability. For a neutral atom, $q_{Z_{0,s}}=0$, so
$\frac{d^{3}\sigma_{Z_{0,s}}^{\text{MY}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}$ $=\bar{q}^{4}\left(\sum_{i}\frac{\overline{A}_{0,s,i}}{\overline{b}_{0,s,i}^{2}+\bar{q}^{2}}\right)^{2}\frac{d^{3}\sigma_{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}},$ (40)
@@
For partially ionized atoms, the complexity is considerably higher because of the presence of the terms $\propto q_{Z_{s^{\prime},s}}\overline{A}_{s^{\prime},s,i}\overline{b}_{s^{\prime},s,i}^{2}$. Let us define
$\overline{c}_{s^{\prime},s,i}^{2}:=q_{Z_{s^{\prime},s}}\overline{b}_{s^{\prime},s,i}^{2},$ (41)
@@
$\frac{d^{3}\sigma_{Z_{s^{\prime},s}}^{\text{MY}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}=\left(\sum_{i}\overline{A}_{s^{\prime},s,i}\frac{\overline{c}_{s^{\prime},s,i}^{2}+\bar{q}^{2}}{\overline{b}_{s^{\prime},s,i}^{2}+\bar{q}^{2}}\right)^{2}\frac{d^{3}\sigma_{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}$ (42)
is obtained leveraging the positivity of $q_{Z_{s^{\prime},s}}$. For the usual Y potential with a single exponential, $i=1$ and since $\overline{A}_{s^{\prime},s,1}=1$ by definition,
$\frac{d^{3}\sigma_{Z_{s^{\prime},s}}^{\text{Y}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}$ $=\frac{d^{3}\sigma_{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}\{\},$ (43)
$\{\}$ $=\overline{c}_{s^{\prime},s}^{4}I_{2,0}\!\left(\overline{b}_{s^{\prime},s},\bar{q}\right)$ (44)
$+2\overline{c}_{s^{\prime},s}^{2}I_{2,1}\!\left(\overline{b}_{s^{\prime},s},\bar{q}\right)+I_{2,2}\!\left(\overline{b}_{s^{\prime},s},\bar{q}\right),$ (45)
@@
$I_{l,m}\left(\overline{b},\bar{q}\right)=\frac{\bar{q}^{2m}}{\left(\overline{b}^{2}+\bar{q}^{2}\right)^{l}}.$ (46)
When screening effects are neglected, $\overline{b}_{s^{\prime},s}=\overline{c}_{s^{\prime},s}=0$, so only the last term in expression (45) remains. In this case,
$\frac{d^{3}\sigma_{Z_{s}}^{\text{Y}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}\Bigg|_{\bar{b}=0}=\frac{d^{3}\sigma^{\text{BH}}_{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}$ (47)
is correctly recovered since $I_{2,2}\left(0,\bar{q}\right)=1$. Conversely, for neutral atoms with screening ($\overline{c}_{0,s}=0$, $\overline{b}_{0,s}\neq 0$) and
$\frac{d^{3}\sigma_{Z_{0,s}}^{\text{Y}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}=I_{2,2}\left(\overline{b}_{0,s},\bar{q}\right)\frac{d^{3}\sigma_{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}},$ (48)
@@
Having established the robustness of the construction through multiple checks, we now derive the cross-section expression for the general MY case. When several exponentials are considered, the screening factor
$\left(\sum_{i}\overline{A}_{s^{\prime},s,i}\frac{\overline{c}_{s^{\prime},s,i}^{2}+\bar{q}^{2}}{\overline{b}_{s^{\prime},s,i}^{2}+\bar{q}^{2}}\right)^{2}$ (49)
@@
$\frac{d^{3}\sigma_{Z_{s^{\prime},s}}^{\text{MY}}}{d\bar{k}\,d\Omega_{k}\,d\Omega_{p}}$ $=\frac{d^{3}\sigma_{Z_{s}}}{d\bar{k}\,d\Omega_{k}\,d\Omega_{p}}$ (50)
$\times\biggl\{\sum_{i}\overline{A}_{s^{\prime},s,i}^{2}\bigl[\overline{c}_{s^{\prime},s,i}^{4}I_{2,0}\left(\overline{b}_{s^{\prime},s,i},\bar{q}\right)+2\overline{c}_{s^{\prime},s,i}^{2}I_{2,1}\left(\overline{b}_{s^{\prime},s,i},\bar{q}\right)+I_{2,2}\left(\overline{b}_{s^{\prime},s,i},\bar{q}\right)\bigr]$ (51)
$\quad+\sum_{i\neq i^{\prime}}\frac{\overline{A}_{s^{\prime},s,i}\,\overline{A}_{s^{\prime},s,i^{\prime}}}{\left(\overline{b}_{s^{\prime},s,i^{\prime}}^{2}-\overline{b}_{s^{\prime},s,i}^{2}\right)}\bigl[\overline{c}_{s^{\prime},s,i}^{2}\overline{c}_{s^{\prime},s,i^{\prime}}^{2}\bigl(I_{1,0}\left(\overline{b}_{s^{\prime},s,i},\bar{q}\right)-I_{1,0}\left(\overline{b}_{s^{\prime},s,i^{\prime}},\bar{q}\right)\bigr)$ (52)
$\quad\quad+\left(\overline{c}_{s^{\prime},s,i}^{2}+\overline{c}_{s^{\prime},s,i^{\prime}}^{2}\right)\bigl(I_{1,1}\left(\overline{b}_{s^{\prime},s,i},\bar{q}\right)-I_{1,1}\left(\overline{b}_{s^{\prime},s,i^{\prime}},\bar{q}\right)\bigr)$ (53)
$\quad\quad+I_{1,2}\left(\overline{b}_{s^{\prime},s,i},\bar{q}\right)-I_{1,2}\left(\overline{b}_{s^{\prime},s,i^{\prime}},\bar{q}\right)\bigr]\biggr\}.$ (54)
The latter expression (54) must be integrated over the scattering directions $\Omega_{p}$ to obtain the screened, ionized doubly differential cross section. It is easily verified that upon setting $\overline{c}_{s^{\prime},s,i}=0$, the neutral case from [17] is recovered and that in the unscreened limit $\overline{b}_{s^{\prime},s,i}^{2}\downarrow 0$, the Bethe–Heitler TDCS is retrieved since $I_{2,2}(0,\bar{q})=1$. Let us define the double integrals $\mathcal{H}_{l,m}$ for $m\in\left\{0,1,2\right\}$ and $l\in\left\{1,2\right\}$,
$\mathcal{H}_{l,m}\left(\overline{b},\bar{q}\right)$ $:=\int_{0}^{\pi}\int_{0}^{2\pi}\frac{d^{3}\sigma_{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}I_{l,m}\left(\overline{b},\bar{q}\right)d\phi\sin\theta d\theta,$
$=\int_{0}^{\pi}\int_{0}^{2\pi}\frac{d^{3}\sigma_{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}\frac{\bar{q}^{2m}}{(\overline{b}^{2}+\bar{q}^{2})^{l}}d\phi\sin\theta d\theta.$ (55)
Note that in order to determine the screened DDCS, only the integrals $\mathcal{H}_{1,m}\left(\overline{b}\right)$ ought to be determined explicitly, since $\mathcal{H}_{2,m}\left(\overline{b}\right)=-\partial\mathcal{H}_{1,m}\left(\overline{b}\right)/\partial\bar{b}^{2}$, as a consequence of the relation
$\frac{\partial I_{1,m}(\overline{b})}{\partial\overline{b}^{2}}=-\frac{\bar{q}^{2m}}{(\overline{b}^{2}+\bar{q}^{2})^{2}}=-I_{2,m}(\overline{b}).$ (56)
@@
$\frac{d^{2}\sigma_{Z_{s^{\prime},s}}^{\text{MY}}}{d\bar{k}d\Omega_{k}}$ $=\sum_{i}\overline{A}_{s^{\prime},s,i}^{2}\left[\overline{c}_{s,s^{\prime},i}^{4}\mathcal{H}_{2,0}\left(\overline{b}_{s^{\prime},s,i}\right)+2\overline{c}_{s^{\prime},s,i}^{2}\mathcal{H}_{2,1}\left(\overline{b}_{s^{\prime},s,i}\right)+\mathcal{H}_{2,2}\left(\overline{b}_{s^{\prime},s,i}\right)\right]$
$+\sum_{i\neq i^{\prime}}\frac{\overline{A}_{s^{\prime},s,i}\overline{A}_{s^{\prime},s,i^{\prime}}}{\left(\overline{b}_{s^{\prime},s,i^{\prime}}^{2}-\overline{b}_{s^{\prime},s,i}^{2}\right)}\left[\overline{c}_{s^{\prime},s,i}^{2}\overline{c}_{s^{\prime},s,i^{\prime}}^{2}\left(\mathcal{H}_{1,0}\left(\overline{b}_{s^{\prime},s,i}\right)-\mathcal{H}_{1,0}\left(\overline{b}_{s^{\prime},s,i^{\prime}}\right)\right)\right$.
$\times\left(\overline{c}_{s^{\prime},s,i}^{2}+\overline{c}_{s^{\prime},s,i^{\prime}}^{2}\right)\left(\mathcal{H}_{1,1}\left(\overline{b}_{s^{\prime},s,i}\right)-\mathcal{H}_{1,1}\left(\overline{b}_{s^{\prime},s,i^{\prime}}\right)\right)$
$+\left.\left(\mathcal{H}_{1,2}\left(\overline{b}_{s^{\prime},s,i}\right)-\mathcal{H}_{1,2}\left(\overline{b}_{s^{\prime},s,i^{\prime}}\right)\right)\right]\!.$ (57)
@@
$\frac{d^{2}\sigma_{Z_{s^{\prime},s}}^{\text{Y}}}{d\bar{k}d\Omega_{k}}=$ $\overline{c}_{s^{\prime},s}^{4}\mathcal{H}_{2,0}\left(\overline{b}_{s^{\prime},s}\right)$ (58)
$+2\overline{c}_{s^{\prime},s}^{2}\mathcal{H}_{2,1}\left(\overline{b}_{s^{\prime},s}\right)+\mathcal{H}_{2,2}\left(\overline{b}_{s^{\prime},s}\right).$ (59)
@@
$\mathcal{H}_{1,0}(\bar{b})$ $=\frac{I_{2}(\bar{b}^{2})-I_{2}(0)+\bar{b}^{2}}{\bar{b}^{4}}$ (60)
$\mathcal{H}_{1,1}(\bar{b})$ $=\frac{I_{2}(0)-I_{2}(\bar{b}^{2})}{\bar{b}^{2}}$ (61)
$\mathcal{H}_{1,2}(\bar{b})$ $=I_{2}(\bar{b})$ (62)
$\mathcal{H}_{2,0}(\bar{b})$ $=2\frac{I_{2}(\bar{b}^{2})-I_{2}(0)+\bar{b}^{2}}{\bar{b}^{6}}+\frac{I_{1}(\bar{b}^{2})-1}{\bar{b}^{4}}$ (63)
$\mathcal{H}_{2,1}(\bar{b})$ $=\frac{I_{2}(0)-I_{2}(\bar{b}^{2})}{\bar{b}^{4}}-\frac{I_{1}(\bar{b}^{2})}{\bar{b}^{2}}$ (64)
$\mathcal{H}_{2,2}(\bar{b})$ $=I_{1}(\bar{b}^{2}).$ (65)
@@
$I_{1}(\bar{b})$ $=\frac{\alpha Z^{2}r_{0}^{2}}{2\pi\,\bar{k}\,\bar{p}_{0}}\Bigg\{16\,\bar{p}\,(4\bar{E}_{0}^{2}+\bar{b}^{2})\frac{\|\bar{\bm{p}}_{0}\times\bar{\bm{k}}\|^{2}}{\bar{k}^{2}W^{4}}$ (66)
$\quad-\frac{2\,\bar{p}}{W^{2}}\Big[\,4\bar{E}_{0}^{2}+2\bar{E}_{0}\bar{E}-\bar{E}\frac{D_{1}}{\bar{k}}+\bar{b}^{2}\big(1-2\bar{E}\frac{\bar{k}}{D_{1}}\big)\,\Big]$ (67)
$\quad+\frac{2\,\bar{p}}{W^{2}}\,\frac{D_{1}^{2}+2(\bar{E}_{0}\bar{E}-1)D_{1}+\bar{b}^{2}(D_{1}-2\bar{E}\bar{k})}{(D_{1}+\bar{b}^{2})^{2}+4\bar{p}^{2}\bar{b}^{2}}$ (68)
$\qquad\times\Bigg[\,\frac{1}{D_{1}}\big(16\bar{E}_{0}\bar{E}-4\bar{E}_{0}^{2}\bar{b}^{2}-b^{4}\big)-\frac{4\bar{E}_{0}^{2}+\bar{b}^{2}}{\bar{k}^{2}W^{2}}\,$ (69)
$\qquad\;\times\{D_{1}^{2}+2(\bar{E}_{0}\bar{E}-1)D_{1}+\bar{b}^{2}(D_{1}-2\bar{E}\bar{k})\}\Bigg]$ (70)
$\quad-\frac{2\bar{k}^{2}\bar{p}}{(D_{1}+\bar{b}^{2})^{2}+4\bar{p}^{2}\bar{b}^{2}}\Bigg[\frac{4}{D_{1}^{2}}\big(4\bar{E}_{0}^{2}+(1-D_{1})\bar{b}^{2}\big)$ (71)
$\qquad+\frac{D_{1}^{2}+2(\bar{E}_{0}\bar{E}-1)D_{1}+\bar{b}^{2}(D_{1}-2\bar{E}\bar{k})}{D_{1}\,\|\,\bar{\bm{p}}_{0}-\bar{\bm{k}}\,\|^{2}}\Bigg]$ (72)
$\quad-\frac{4\bar{k}}{D_{1}}\ln(\bar{E}+\bar{p})+\frac{L_{1}}{W}\Bigg[\,2\bar{k}+\frac{4\bar{k}}{D_{1}}(\bar{E}_{0}^{2}+\bar{p}^{2}+\bar{b}^{2})\,$ (73)
$\quad+\frac{2}{D_{1}W^{2}}(\bar{E}_{0}D_{1}-2\bar{k}+\bar{b}^{2}\bar{k})$ (74)
$\qquad\times\Big(8\bar{E}_{0}\bar{E}-\frac{D_{1}^{2}}{2}-\bar{b}^{2}(2\bar{E}_{0}^{2}+2\bar{p}^{2}+D_{1})-b^{4}\Big)$ (75)
$\quad+\frac{2(2\bar{E}_{0}^{2}+\bar{b}^{2})(D_{1}-2\bar{E}\bar{k})+D_{1}^{2}+2(\bar{E}_{0}\bar{E}-1)D_{1}}{\bar{k}\,W^{2}}$ (76)
$\quad-\frac{3}{\bar{k}\,W^{4}}(4\bar{E}_{0}^{2}+\bar{b}^{2})\Big(\bar{E}_{0}\frac{D_{1}}{\bar{k}}-2+\bar{b}^{2}\Big)$ (77)
$\qquad\big(D_{1}^{2}+2(\bar{E}_{0}\bar{E}-1)D_{1}+\bar{b}^{2}(D_{1}-2\bar{E}\bar{k})\big)\Bigg]$ (78)
$+\frac{\bar{k}^{2}L_{2}}{D_{1}\,\|\,\bar{\bm{p}}_{0}-\bar{\bm{k}}\,\|}\Big[\,\frac{2}{D_{1}}-2+\frac{D_{1}-2\bar{E}\bar{k}}{2\,\|\,\bar{\bm{p}}_{0}-\bar{\bm{k}}\,\|^{2}}\,\Big]\Bigg\}\,$, (79)

### v2

@@ III.2 Screened Cross Section
$\frac{d^{3}\sigma_{Z_{s}}^{\text{MY}}}{d{k}d\Omega_{k}d\Omega_{p}}=\alpha\left(\frac{r_{e}}{2\pi}\right)^{2}Z^{2}\times$ (30)
$\left(1-\left(1-q_{Z_{s}}\right)\sum_{i}\frac{{A}_{s,i}}{1+\left({q}{a}_{Z_{s,i}}/2\right)^{2}}\right)^{2}\frac{{p}}{{k}{p}_{0}{q}^{4}}A^{\text{BH}},$ (31)
where the term $q_{Z_{s}}:=Z_{s}/Z$, ranging between 0 and $1$, characterizes the level of ionization, 0 being the neutral atom. Defining $b_{s,i}:=2/{a}_{Z_{s,i}}$, Eq. (30) may then be rewritten in the form
$\frac{d^{3}\sigma_{Z_{s}}^{\text{MY}}}{d{k}d\Omega_{k}d\Omega_{p}}=\left(\sum_{i}{A}_{s,i}\frac{{q}^{2}+q_{Z_{s}}{b}_{s,i}^{2}}{{q}^{2}+{b}_{s,i}^{2}}\right)^{2}\frac{d^{3}\sigma_{Z}^{\text{BH}}}{d{k}d\Omega_{k}d\Omega_{p}}.$ (32)
From now on, we omit the BH superscript on the Bethe–Heitler TDCS for the sake of readability. For a neutral atom, $q_{Z_{0}}=0$, so
$\frac{d^{3}\sigma_{Z_{0}}^{\text{MY}}}{d{k}d\Omega_{k}d\Omega_{p}}$ $={q}^{4}\left(\sum_{i}\frac{{A}_{0,i}}{{b}_{0,i}^{2}+{q}^{2}}\right)^{2}\frac{d^{3}\sigma_{Z}}{d{k}d\Omega_{k}d\Omega_{p}},$ (33)
@@
For partially ionized atoms, the complexity is considerably higher because of the presence of the terms $\propto q_{Z_{s}}{A}_{s,i}{b}_{s,i}^{2}$. Let us define
${c}_{s,i}^{2}:=q_{Z_{s}}{b}_{s,i}^{2},$ (34)
@@
$\frac{d^{3}\sigma_{Z_{s}}^{\text{MY}}}{d{k}d\Omega_{k}d\Omega_{p}}=\left(\sum_{i}{A}_{s,i}\frac{{c}_{s,i}^{2}+{q}^{2}}{{b}_{s,i}^{2}+{q}^{2}}\right)^{2}\frac{d^{3}\sigma_{Z}}{d{k}d\Omega_{k}d\Omega_{p}}$ (35)
is obtained leveraging the positivity of $q_{Z_{s}}$. For the usual Y potential with a single exponential, $i=1$ and since ${A}_{s,1}=1$ by definition,
$\frac{d^{3}\sigma_{Z_{s}}^{\text{Y}}}{d{k}d\Omega_{k}d\Omega_{p}}$ $=\frac{d^{3}\sigma_{Z}}{d{k}d\Omega_{k}d\Omega_{p}}\{\},$ (36)
$\{\}$ $={c}_{s}^{4}I_{2,0}\!\left({b}_{s},{q}\right)$ (37)
$+2{c}_{s}^{2}I_{2,1}\!\left({b}_{s},{q}\right)+I_{2,2}\!\left({b}_{s},{q}\right),$ (38)
@@
$I_{l,m}\left({b},{q}\right)=\frac{{q}^{2m}}{\left({b}^{2}+{q}^{2}\right)^{l}}.$ (39)
When screening effects are neglected, $b_{s}=c_{s}=0$, so only the last term in expression (38) remains. In this case, the Bethe–Heitler TDCS is correctly recovered since $I_{2,2}\left(0,q\right)=1$. Conversely, for neutral atoms with screening ($c_{0,s}=0$, $b_{0,s}\neq 0$) and
$\frac{d^{3}\sigma_{Z_{0}}^{\text{Y}}}{d{k}d\Omega_{k}d\Omega_{p}}=I_{2,2}\left({b}_{0},{q}\right)\frac{d^{3}\sigma_{Z}}{d{k}d\Omega_{k}d\Omega_{p}},$ (40)
@@
Having established the robustness of the construction through multiple checks, we now derive the doubly-differential cross section for the general MY case. When several exponentials are considered, the screening factor
$\left(\sum_{i}{A}_{s,i}\frac{{c}_{s,i}^{2}+{q}^{2}}{{b}_{s,i}^{2}+{q}^{2}}\right)^{2}$ (41)
@@
$\frac{d^{3}\sigma_{Z_{s}}^{\text{MY}}}{d{k}\,d\Omega_{k}\,d\Omega_{p}}$ $=\frac{d^{3}\sigma_{Z}}{d{k}\,d\Omega_{k}\,d\Omega_{p}}$ (42)
$\times\biggl\{\sum_{i}{A}_{s,i}^{2}\bigl[{c}_{s,i}^{4}I_{2,0}\left({b}_{s,i},{q}\right)+2{c}_{s,i}^{2}I_{2,1}\left({b}_{s,i},{q}\right)+I_{2,2}\left({b}_{s,i},{q}\right)\bigr]$ (43)
$\quad+\sum_{i\neq i^{\prime}}\frac{{A}_{s,i}\,{A}_{s,i^{\prime}}}{\left({b}_{s,i^{\prime}}^{2}-{b}_{s,i}^{2}\right)}\bigl[{c}_{s,i}^{2}{c}_{s,i^{\prime}}^{2}\bigl(I_{1,0}\left({b}_{s,i},{q}\right)-I_{1,0}\left({b}_{s,i^{\prime}},{q}\right)\bigr)$ (44)
$\quad\quad+\left({c}_{s,i}^{2}+{c}_{s,i^{\prime}}^{2}\right)\bigl(I_{1,1}\left({b}_{s,i},{q}\right)-I_{1,1}\left({b}_{s,i^{\prime}},{q}\right)\bigr)$ (45)
$\quad\quad+I_{1,2}\left({b}_{s,i},{q}\right)-I_{1,2}\left({b}_{s,i^{\prime}},{q}\right)\bigr]\biggr\}.$ (46)
The latter expression (46) must be integrated over the scattering directions $\Omega_{p}$ to obtain the screened, ionized doubly differential cross section. It is easily verified that upon setting $c_{s,i}=0$, the neutral case from [18] is recovered and that in the unscreened limit ${b}_{s,i}^{2}\downarrow 0$, the Bethe–Heitler TDCS is retrieved. Let us define the double integrals $\mathcal{H}_{l,m}$ for $m\in\left\{0,1,2\right\}$ and $l\in\left\{1,2\right\}$,
$\mathcal{H}_{l,m}\left({b},{q}\right)$ $:=\int_{0}^{\pi}\int_{0}^{2\pi}\frac{d^{3}\sigma_{Z}}{d{k}d\Omega_{k}d\Omega_{p}}I_{l,m}\left({b},{q}\right)d\phi\sin\theta d\theta,$
$=\int_{0}^{\pi}\int_{0}^{2\pi}\frac{d^{3}\sigma_{Z}}{d{k}d\Omega_{k}d\Omega_{p}}\frac{{q}^{2m}}{({b}^{2}+{q}^{2})^{l}}d\phi\sin\theta d\theta.$ (47)
Note that in order to determine the screened DDCS, only the integrals $\mathcal{H}_{1,m}\left({b}\right)$ ought to be determined explicitly, since $\mathcal{H}_{2,m}\left({b}\right)=-\partial\mathcal{H}_{1,m}\left({b}\right)/\partial{b}^{2}$, as a consequence of the relation
$\frac{\partial I_{1,m}({b})}{\partial{b}^{2}}=-\frac{{q}^{2m}}{({b}^{2}+{q}^{2})^{2}}=-I_{2,m}({b}).$ (48)
@@
$\frac{d^{2}\sigma_{Z_{s}}^{\text{MY}}}{d{k}d\Omega_{k}}$ $=\sum_{i}{A}_{s,i}^{2}\left[{c}_{s,i}^{4}\mathcal{H}_{2,0}\left({b}_{s,i}\right)+2{c}_{s,i}^{2}\mathcal{H}_{2,1}\left({b}_{s,i}\right)+\mathcal{H}_{2,2}\left({b}_{s,i}\right)\right]$
$+\sum_{i\neq i^{\prime}}\frac{{A}_{s,i}{A}_{s,i^{\prime}}}{\left({b}_{s,i^{\prime}}^{2}-{b}_{s,i}^{2}\right)}\left[{c}_{s,i}^{2}{c}_{s,i^{\prime}}^{2}\left(\mathcal{H}_{1,0}\left({b}_{s,i}\right)-\mathcal{H}_{1,0}\left({b}_{s,i^{\prime}}\right)\right)\right$.
$+\left({c}_{s,i}^{2}+{c}_{s,i^{\prime}}^{2}\right)\left(\mathcal{H}_{1,1}\left({b}_{s,i}\right)-\mathcal{H}_{1,1}\left({b}_{s,i^{\prime}}\right)\right)$
$+\left.\left(\mathcal{H}_{1,2}\left({b}_{s,i}\right)-\mathcal{H}_{1,2}\left({b}_{s,i^{\prime}}\right)\right)\right]\!.$ (49)
@@
$\frac{d^{2}\sigma_{Z_{s}}^{\text{Y}}}{d{k}d\Omega_{k}}=$ ${c}_{s}^{4}\mathcal{H}_{2,0}\left({b}_{s}\right)$ (50)
$+2{c}_{s}^{2}\mathcal{H}_{2,1}\left({b}_{s}\right)+\mathcal{H}_{2,2}\left({b}_{s}\right).$ (51)
@@
$\mathcal{H}_{1,0}({b})$ $=\frac{I_{2}({b}^{2})-I_{2}(0)+{b}^{2}}{{b}^{4}}$ (52)
$\mathcal{H}_{1,1}({b})$ $=\frac{I_{2}(0)-I_{2}({b}^{2})}{{b}^{2}}$ (53)
$\mathcal{H}_{1,2}({b})$ $=I_{2}({b})$ (54)
$\mathcal{H}_{2,0}({b})$ $=2\frac{I_{2}({b}^{2})-I_{2}(0)+{b}^{2}}{{b}^{6}}+\frac{I_{1}({b}^{2})-1}{{b}^{4}}$ (55)
$\mathcal{H}_{2,1}({b})$ $=\frac{I_{2}(0)-I_{2}({b}^{2})}{{b}^{4}}-\frac{I_{1}({b}^{2})}{{b}^{2}}$ (56)
$\mathcal{H}_{2,2}({b})$ $=I_{1}({b}^{2}).$ (57)
@@
$I_{1}({b})$ $=\frac{\alpha Z^{2}r_{0}^{2}}{2\pi\,{k}\,{p}_{0}}\Bigg\{16\,{p}\,(4{E}_{0}^{2}+{b}^{2})\frac{\|{\bm{p}}_{0}\times{\bm{k}}\|^{2}}{{k}^{2}W^{4}}$ (58)
$\quad-\frac{2\,{p}}{W^{2}}\Big[\,4{E}_{0}^{2}+2{E}_{0}{E}-{E}\frac{D_{1}}{{k}}+{b}^{2}\big(1-2{E}\frac{{k}}{D_{1}}\big)\,\Big]$ (59)
$\quad+\frac{2\,{p}}{W^{2}}\,\frac{D_{1}^{2}+2({E}_{0}{E}-1)D_{1}+{b}^{2}(D_{1}-2{E}{k})}{(D_{1}+{b}^{2})^{2}+4{p}^{2}{b}^{2}}$ (60)
$\qquad\times\Bigg[\,\frac{1}{D_{1}}\big(16{E}_{0}{E}-4{E}_{0}^{2}{b}^{2}-b^{4}\big)-\frac{4{E}_{0}^{2}+{b}^{2}}{{k}^{2}W^{2}}\,$ (61)
$\qquad\;\times\{D_{1}^{2}+2({E}_{0}{E}-1)D_{1}+{b}^{2}(D_{1}-2{E}{k})\}\Bigg]$ (62)
$\quad-\frac{2{k}^{2}{p}}{(D_{1}+{b}^{2})^{2}+4{p}^{2}{b}^{2}}\Bigg[\frac{4}{D_{1}^{2}}\big(4{E}_{0}^{2}+(1-D_{1}){b}^{2}\big)$ (63)
$\qquad+\frac{D_{1}^{2}+2({E}_{0}{E}-1)D_{1}+{b}^{2}(D_{1}-2{E}{k})}{D_{1}\,\|\,{\bm{p}}_{0}-{\bm{k}}\,\|^{2}}\Bigg]$ (64)
$\quad-\frac{4{k}}{D_{1}}\ln({E}+{p})+\frac{L_{1}}{W}\Bigg[\,2{k}+\frac{4{k}}{D_{1}}({E}_{0}^{2}+{p}^{2}+{b}^{2})\,$ (65)
$\quad+\frac{2}{D_{1}W^{2}}({E}_{0}D_{1}-2{k}+{b}^{2}{k})$ (66)
$\qquad\times\Big(8{E}_{0}{E}-\frac{D_{1}^{2}}{2}-{b}^{2}(2{E}_{0}^{2}+2{p}^{2}+D_{1})-b^{4}\Big)$ (67)
$\quad+\frac{2(2{E}_{0}^{2}+{b}^{2})(D_{1}-2{E}{k})+D_{1}^{2}+2({E}_{0}{E}-1)D_{1}}{{k}\,W^{2}}$ (68)
$\quad-\frac{3}{{k}\,W^{4}}(4{E}_{0}^{2}+{b}^{2})\Big({E}_{0}\frac{D_{1}}{{k}}-2+{b}^{2}\Big)$ (69)
$\qquad\big(D_{1}^{2}+2({E}_{0}{E}-1)D_{1}+{b}^{2}(D_{1}-2{E}{k})\big)\Bigg]$ (70)
$+\frac{{k}^{2}L_{2}}{D_{1}\,\|\,{\bm{p}}_{0}-{\bm{k}}\,\|}\Big[\,\frac{2}{D_{1}}-2+\frac{D_{1}-2{E}{k}}{2\,\|\,{\bm{p}}_{0}-{\bm{k}}\,\|^{2}}\,\Big]\Bigg\}\,$, (71)